# Motion Matters: Motion-guided Modulation Network for Skeleton-based Micro-Action Recognition

### v2

@@ 1. Introduction
In human-centric micro-action analysis (Liu et al., 2021; Chen et al., 2023; Balazia et al., 2022; Guo et al., 2024a; Li et al., 2023c, b), remarkable progress has been made. From the perspective of the dataset, Liu et al. and Chen et al. collected the iMiGUE (Liu et al., 2021) and SMG (Chen et al., 2023) datasets that contain spontaneous micro-gestures in the upper limbs. These videos are sampled from the post-competition review of athletes, revealing deep emotional states conveyed through these micro-gestures. Meanwhile, Balazia et al. presented the MPIIGI (Balazia et al., 2022) dataset, which annotated the subtle upper-body behaviors in group interactions and highlighted the social behaviors in human communications. To explore the whole body micro-actions, Guo et al(Guo et al., 2024a) developed the large-scale Micro-Action 52 (MA-52) dataset comprising 20k video samples interviewed from 205 participants, distributed in 52 micro-action classes, benchmarking micro-action recognition. From the perspective of methodology, traditional action recognition methods, including 2D CNN based (Wang et al., 2018; Lin et al., 2019; Feichtenhofer et al., 2019), 3D CNN based (Tran et al., 2015; Carreira and Zisserman, 2017), GCN-based (Yan et al., 2018; Liu et al., 2020), and Transformer-based (Liu et al., 2022), are selected to evaluate these datasets. However, these methods show suboptimal performance due to the challenges posed by the subtle and low-intensity nature of micro-actions. Recently, Li et al(Li et al., 2025a) proposed a novel prototypical ambiguous calibrating net to address the ambiguity in micro-actions.
@@
  • We propose a skeletal-temporal context-aware augmentation strategy to enhance the robustness and generalization of the proposed method in the context of micro-action recognition.
@@ 2.1. Micro-Action Recognition
2.2. Skeleton-based Action Recognition
@@ 3.3. Feature Embedding
where $t$, $v$, and $cs$ denote the temporal step, joint index, and channel index, respectively.
@@ 3.4. Motion-guided Feature Modulation
To capture the subtle motion patterns in the skeleton sequence, we propose a Motion-guided Feature Modulation (MFM) module that models complex motion dynamics to guide the learning of compact spatio-temporal features. As shown in Figure 2, MFM is composed of $N$ stacked Motion-guided Skeletal-Temporal Formers (MSTF). Inspired by Transformer-based methods (Wang and Koniusz, 2023; Do and Kim, 2024) in skeleton-based action recognition, we utilize three parallel branches—namely, the skeletal branch, motion branch, and temporal branch—to construct compact spatio-temporal representations. The input is first normalized and projected as:
@@
where $\cdot$] denotes channel-wise concatenation, $\texttt{Aggregation}(\cdot)$ learns different weights based on the feature distributions of each branch. This strategy helps balance contributions from spatial, temporal, and motion-specific features. To further enhance network stability, the fused features ${\bm{X}}_{agg}$ pass through a projection layer and are then added via a residual connection to the original input ${\bm{X}}_{raw}$. A FeedForward network (Vaswani et al., 2017) is subsequently applied to introduce nonlinear transformations. Overall, the stacked $N$ layers Motion-guided Skeletal-Temporal Former preserve deeper motion cues, strengthening the final representations ${\bm{X}}_{f}$ of micro-action movements.
@@ 3.5. Motion Consistency Learning
Micro-actions span a wide range of categories, specifically, 52 action-level categories derived from 7 body-level categories (e.g., head, hand, and body). As a result, different micro-actions exhibit variations in temporal duration. To capture motion patterns across multiple temporal scales, we construct $L$ spatio-temporal features (${\bm{X}}_{f}^{1},{\bm{X}}_{f}^{2},\ldots,{\bm{X}}_{f}^{L}$) by progressively downsampling the temporal dimension by a factor of two at each stage. Then, each scale feature is input to the Motion-guided Feature Modulation to build coarse to fine spatio-temporal features. Specifically, the unified temporal length is denoted as $T^{\prime}=T/2^{(L-1)}$. Subsequently, we utilize an adaptive gated fusion module to aggregate multi-scale representations:
@@ 4.1. Experimental Setup
Datasets. Micro-Action 52 (Guo et al., 2024a) is a large-scale whole-body micro-action dataset collected by a professional interview to capture unconscious human micro-action behaviors. The dataset contains 22,422 (22.4K) samples, where the annotations are categorized into two levels: 7 body-level and 52 action-level micro-action categories. There are 11,250, 5,586, and 5,586 instances in the training, validation, and test sets, respectively. In addition, we also evaluate the proposed method on a micro-action-related dataset, i.e., iMiGUE (Liu et al., 2021), which mainly consists of micro-gestures of the upper body. iMiGUE was collected from post-competition athlete interviews and contains 32 micro-gesture categories. While conceptually similar to micro-actions, these micro-gestures are restricted to the upper limbs. Due to privacy concerns, we can only get the skeleton data released by the MiGA Competition (Chen et al., 2024a). Therefore, we only report the results of skeleton-based methods. The dataset contains 12,893, 777, and 4,562 samples for training, validation, and testing, respectively.
@@ 4.3. Ablation Studies
Skeletal Temporal Top-1 Top-1 Top-5 F1mean

### v3

@@ 1. Introduction
In human-centric micro-action analysis (Liu et al., 2021; Chen et al., 2023; Balazia et al., 2022; Guo et al., 2024a; Li et al., 2023c, b), remarkable progress has been made. From the perspective of the dataset, Liu et al. and Chen et al. collected the iMiGUE (Liu et al., 2021) and SMG (Chen et al., 2023) datasets that contain spontaneous micro-gestures in the upper limbs. These videos are sampled from the post-competition review of athletes, revealing deep emotional states conveyed through these micro-gestures. Meanwhile, Balazia et al(Balazia et al., 2022) presented the MPIIGI dataset, which annotated the subtle upper-body behaviors in group interactions and highlighted the social behaviors in human communications. To explore the whole body micro-actions, Guo et al(Guo et al., 2024a) developed the large-scale Micro-Action 52 (MA-52) dataset comprising 20k video samples interviewed from 205 participants, distributed in 52 micro-action classes, benchmarking micro-action recognition. From the perspective of methodology, traditional action recognition methods, including 2D CNN based (Wang et al., 2018; Lin et al., 2019; Feichtenhofer et al., 2019), 3D CNN based (Tran et al., 2015; Carreira and Zisserman, 2017), GCN-based (Yan et al., 2018; Liu et al., 2020), and Transformer-based (Liu et al., 2022), are selected to evaluate these datasets. However, these methods show suboptimal performance due to the challenges posed by the subtle and low-intensity nature of micro-actions. Recently, Li et al(Li et al., 2025a) proposed a novel prototypical ambiguous calibrating net to address the ambiguity in micro-actions.
@@
  • We conduct a comprehensive evaluation of existing skeleton-based action recognition methods in the context of micro-action recognition and establish standardized protocols for benchmark datasets.
@@ 2.1. Micro-Action Recognition
$2.$ Skeleton-based Action Recognition
@@ 3.2. Network Overview
where $t$, $v$, and cscs denote the temporal step, joint index, and channel index, respectively.
@@ 3.4. Motion-guided Feature Modulation
To capture the subtle motion patterns in the skeleton sequence, we propose a Motion-guided Feature Modulation (MFM) module that models complex motion dynamics to guide the learning of compact spatio-temporal features. As shown in Figure 2, MFM is composed of NN stacked Motion-guided Skeletal-Temporal Formers (MSTF). Inspired by Transformer-based methods (Wang and Koniusz, 2023; Do and Kim, 2024) in skeleton-based action recognition, we utilize three parallel branches—namely, the skeletal branch, motion branch, and temporal branch—to construct compact spatio-temporal representations. The input is first normalized and projected as:
@@
where $\cdot$] denotes channel-wise concatenation, $\texttt{Aggregation}(\cdot)$ learns different weights based on the feature distributions of each branch. This strategy helps balance contributions from spatial, temporal, and motion-specific features. To further enhance network stability, the fused features ${\bm{X}}_{agg}$ pass through a projection layer and are then added via a residual connection to the original input ${\bm{X}}_{raw}$. A FeedForward network (Vaswani et al., 2017) is subsequently applied to introduce nonlinear transformations. Overall, the stacked NN-layer Motion-guided Skeletal-Temporal Former preserves deeper motion cues, thereby strengthening the final representations ${\bm{X}}_{f}$ of micro-action movements.
@@ 3.5. Motion Consistency Learning
Micro-actions span a wide range of categories, specifically, 52 action-level categories derived from 7 body-level categories (e.g., head, hand, and body). As a result, different micro-actions exhibit variations in temporal duration. To capture motion patterns across multiple temporal scales, we construct $L$ spatio-temporal features (${\bm{X}}_{f}^{1},{\bm{X}}_{f}^{2},\ldots,{\bm{X}}_{f}^{L}$) by progressively downsampling the temporal dimension by a factor of two at each stage. Then, each scale feature is input to the Motion-guided Feature Modulation to build coarse-to-fine spatio-temporal features. Specifically, the unified temporal length is denoted as $T^{\prime}=T/2^{(L-1)}$. Subsequently, we utilize an adaptive gated fusion module to aggregate multi-scale representations:
@@ 4.1. Experimental Setup
Datasets. Micro-Action 52 (Guo et al., 2024a) is a large-scale whole-body micro-action dataset collected by a professional interview to capture unconscious human micro-action behaviors. The dataset contains 22,422 (22.4K) samples, where the annotations are categorized into two levels: 7 body-level and 52 action-level micro-action categories. There are 11,250, 5,586, and 5,586 instances in the training, validation, and test sets, respectively. In addition, we also evaluate the proposed method on a micro-action-related dataset, i.e., iMiGUE (Liu et al., 2021), which mainly consists of micro-gestures of the upper body. iMiGUE was collected from post-competition athlete interviews and contains 32 micro-gesture categories. While conceptually similar to micro-actions, these micro-gestures are restricted to the upper limbs. We use the skeleton data released by the MiGA Competition (Chen et al., 2024a). The dataset contains 12,893, 777, and 4,562 samples for training, validation, and testing, respectively.
@@ 4.3. Ablation Studies
Exp. Skeletal Temporal Top-1 Top-1 Top-5 F1mean